# Self-Correcting Code Generation Using Small Language Models

### v1

@@ Abstract
Self-correction has demonstrated potential in code generation by allowing language models to revise and improve their outputs through successive refinement. Recent studies have explored prompting-based strategies that incorporate verification or feedback loops using proprietary models, as well as training-based methods that leverage their strong reasoning capabilities. However, whether smaller models possess the capacity to effectively guide their outputs through self-reflection remains unexplored. Our findings reveal that smaller models struggle to exhibit reflective revision behavior across both self-correction paradigms. In response, we introduce CoCoS, an approach designed to enhance the ability of small language models for multi-turn code correction. Specifically, we propose an online reinforcement learning objective that trains the model to confidently maintain correct outputs while progressively correcting incorrect outputs as turns proceed. Our approach features an accumulated reward function that aggregates rewards across the entire trajectory and a fine-grained reward better suited to multi-turn correction scenarios. This facilitates the model in enhancing initial response quality while achieving substantial improvements through self-correction. With 1B-scale models, CoCoS achieves improvements of 35.8% on the MBPP and 27.7% on HumanEval compared to the baselines.111Our implementation can be accessed at https://github.com/jeonghun3572/CoCoS.
@@
Jeonghun Cho1, Deokhyung Kang1, Hyounghun Kim1,2, Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence 2Department of Computer Science and Engineering Pohang University of Science and Technology (POSTECH) {jeonghuncho, deokhk, h.kim, gblee}@postech.ac.kr
@@ 3 Problem Description
In this work, we focus on code generation tasks and study a small-scale LLM to investigate its intrinsic ability to improve its outputs over subsequent trials. Suppose consists of data points, each represented as a tuple , where is a problem, is the canonical code snippet, and is the unit test cases. We define the SLM as and aim to generate correct through the conditional distribution , where represents the previous model trials, and denotes auxiliary instructions for revising the previous responses. Note that the subscripts indicate turn indices within the multi-turn setting. Since we do not consider external feedback, is fixed for each turn.
@@ 4.2 Setup
We experiment with two prompting-based self-correction methods, Self-Refine (Madaan et al., 2023) and SETS (Chen et al., 2025)—both demonstrated exclusively on proprietary LLMs in their original work. To verify the effectiveness of SLMs, all experiments are conducted using the pre-trained Qwen2.5-1.5B (Yang et al., 2024). Self-Refine uses 3-shot prompting, whereas SETS relies on test-time scaling. Since SETS does not generate a fixed initial response, we approximate Accuracy@t1 by averaging the performance of the sampled initial outputs. Also, since SETS does not revise verified responses, detailed correction metrics are unobservable. Instead, we report feedback accuracy in Appendix A.
We further investigate the influence of external feedback by employing strong teacher models,222We use gpt-4.1-2025-04-14 (OpenAI, 2025) for teacher model. which we refer to as External-Refine and External-SETS. For comparison, we also introduce a simple correction prompting method that first generates an initial response with a 3-shot prompt, then prompts for correction without disclosing the correctness of the initial output. Additional details including prompts and hyperparameters are provided in Appendix F.
@@ 5.1 Multi-turn MDP
As outlined in §3, our goal is to enhance model outputs across multiple turns by leveraging the response history. Since the model’s own generated history naturally defines a sequential decision process, we formulate the process of self-correction as a multi-turn Markov decision process (Qu et al., 2024), where the policy is defined as . In this formulation, the state at each turn consists of the prior outputs . Given the prompt history , the model generates a revised output , which is then appended to form the next prompt . We evaluate the correctness of the generated code using unit test execution , which assesses the functional correctness of the code against the given unit tests. The test outcome is used as the reward:
where denotes the reward for . Under this formulation, the learning objective is to optimize a policy that maximizes the expected cumulative (with in our setup) reward over a refinement trajectory:
@@ 5.2 Observed limitations
where is a binary reward, taking a value of 1 for all-pass and 0 for fail. The second KL-divergence term represents the regularization, and . However, we find that introducing the second KL-divergence term () leads to training collapse for SLMs, as shown in Figure 1 where the reward converges to zero.
@@ 5.3 CoCoS
Since our setting is also formulated as a multi-turn MDP, we adopt a REINFORCE-style policy gradient method (Ahmadian et al., 2024) to optimize the policy, which is widely used in single-turn RLHF (). In order to adapt to a multi-turn setting, we introduce two key modifications to the reward function. First, we introduce an accumulated reward function that incorporates a discount factor to encourage the model to consider the assessment of the previous turn ( in the case of SCoRe). Unlike conventional RL, where discounts future rewards, we adapt its usage to amplify the contribution of the most recent turn. The accumulated reward function is defined as follows:
@@
The effectiveness of the progressive reward compared to the binary scheme is presented in §8.2. Accordingly, our final training objective is defined as follows:
where is the KL coefficient. Detailed hyperparameters are provided in Appendix C.
| Method | Accuracy@t1 | Accuracy@t2 | (t1, t2) | (t1, t2) | (t1, t2) |
| Boost model | 17.6% | 18.0% | 0.8% | 0.4% | 0.4% |
| Turn-SFT | 44.2% | 44.4% | 1.4% | 1.2% | 0.2% |
| Self-Corrector | 43.8% | 48.8% | 11.2% | 6.2% | 5.0% |
| ReVISE | 34.6% | 41.2% | 11.3% | 4.9% | 6.4% |
| Turn-RL | 46.4% | 48.6% | 4.0% | 1.8% | 2.2% |
[…]
@@ 6.2 Datasets
While supervised fine-tuning (SFT) allows models to learn proper code formatting through input-output pairs, online-RL lacks such structural guidance. Therefore, we perform SFT on the KodCode (Xu et al., 2025) dataset to produce parseable code so that RL training can be conducted with unit test execution. KodCode is curated to exclude data from MBPP, HumanEval, and ODEX. We measure the model’s accuracy on MBPP at both the pre-SFT and post-SFT, which we denote as the Pre-trained and Boost model, respectively, and report the results in Table 5.3. As shown in Table 5.3, we emphasize that the purpose of SFT on KodCode is not to enhance performance on the target dataset. The Boost model is the backbone for both CoCoS and all baseline comparisons.
@@ 7 Main Results
As shown in Table 5.3, CoCoS demonstrates the highest correction performance across all turns. Although the baselines are trained to actively revise prior responses, the metric indicates that they frequently make unnecessary revisions to already correct outputs. CoCoS, on the other hand, corrects selectively: it achieves a low rate of 1.8%, avoiding unnecessary changes, while maintaining a high rate of 11%. This results in a of 9.2%, the highest among all baselines.
@@ 8 Analyses
In this section, we analyze the discount factor in the accumulated reward function (§8.1), compare progressive and binary reward schemes (§8.2), examine multi-turn correction beyond two turns (§8.3), and evaluate the model’s robustness to varied correction instructions during inference (§8.4).
@@ 8.4 Auxiliary instruction
To evaluate the generalization ability of our method across different prompts, we intentionally use a different instruction prompt at test time than the one used during training (as defined in Appendix F.1). Instead of the fixed training prompt, we manually rephrase it into five alternative versions for the correction task. The results are shown in Table 6, and the rephrased instructions are provided in Appendix G. While the fixed prompt used during training yields slightly better overall performance, the model exhibits no substantial degradation under instruction variation. The average performance drop at the second attempt is only 0.8%, indicating robustness to diverse prompt variations.
@@ 9 Conclusion
Online reinforcement learning incurs a cost at every action step, as the model learns by interacting with the environment in real-time. Due to infrastructural constraints, we limited CoCoS training to only two turns. In the turn-wise analysis section, CoCoS consistently demonstrated accuracy gains as the number of turns increased. These results suggest that increasing the number of training turns has the potential to yield further improvements. Therefore, future work should extend our approach to support multi-turn training beyond two turns, potentially incorporating more cost-efficient strategies such as offline reinforcement learning. For the same reason, our experiments were conducted only on a 1B-scale small language model. While we demonstrated that our methodologies are effective on small models, our approach is not limited to this setting. Moreover, evaluating its scalability to models larger than 1B parameters remains an interesting aspect to observe in future work.
@@
Acknowledgments
This work was supported by the IITP(Institute of Information & Coummunications Technology Planning & Evaluation)-ITRC(Information Technology Research Center) grant funded by the Korea government(Ministry of Science and ICT)(IITP-2025-RS-2024-00437866, 47.5%).
This work was supported by Smart HealthCare Program funded by the Korean National Police Agency(KNPA) (No. RS-2022-PT000186, 47.5%).
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH), 5%).
Appendix A Detailed Evaluation of SETS
@@ Appendix B Baseline Details
We generate training data using pre-trained models according to the format specified in §F.1. For each problem, we sample 10 candidate code solutions and determine their correctness based on the results of corresponding unit test cases. We then organize this data into two separate datasets: 1-turn and 2-turn. We combine the 1-turn and 2-turn datasets in a 1:1 ratio to match the data scale of other baseline methods, allowing the model to optimize initial code generation and correction capabilities.
@@ Appendix C Experimental Details
C.1 Implementation
@@
| Discount factor () | - | 0.5 |
| RLOO samples | - | 2 |
C.2 Dataset
@@ Appendix E Additional SCoRe Experiment on Another Model
Appendix F Prompts
In this section, we present the prompts used for prompt-based approaches and for CoCoS training and evaluation. The prompts for implementing prompting-based approaches are described in §F.2 and §F.3. Finally, the prompts used for CoCoS training and evaluation are reported in §F.1.
F.1 Main experiments

### v3

@@ Abstract
Self-correction has demonstrated potential in code generation by allowing language models to revise and improve their outputs through successive refinement. Recent studies have explored prompting-based strategies that incorporate verification or feedback loops using proprietary models, as well as training-based methods that leverage their strong reasoning capabilities. However, whether smaller models possess the capacity to effectively guide their outputs through self-reflection remains unexplored. Our findings reveal that smaller models struggle to exhibit reflective revision behavior across both self-correction paradigms. In response, we introduce CoCoS, an approach designed to enhance the ability of small language models for multi-turn code correction. Specifically, we propose an online reinforcement learning objective that trains the model to confidently maintain correct outputs while progressively correcting incorrect outputs as turns proceed. Our approach features an accumulated reward function that aggregates rewards across the entire trajectory and a fine-grained reward better suited to multi-turn correction scenarios. This facilitates the model in enhancing initial response quality while achieving substantial improvements through self-correction. With 1B-scale models, CoCoS achieves improvements of 35.8% on the MBPP and 27.7% on HumanEval compared to the baselines.111Our implementation can be accessed at https://github.com/jeonghun3572/CoCoS
@@
Jeonghun Cho1, Deokhyung Kang1, Hyounghun Kim1,2, Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence, POSTECH 2Department of Computer Science and Engineering, POSTECH {jeonghuncho, deokhk, h.kim, gblee}@postech.ac.kr
@@ 3 Problem Description
In this work, we focus on code generation tasks and study a small-scale LLM to investigate its intrinsic ability to improve its outputs over subsequent trials. Suppose consists of data points, each represented as a tuple , where is a problem, is the canonical code snippet, and is the unit test cases. We define the SLM as and aim to generate correct through the conditional distribution , where represents the history of SLM’s previous trials, and denotes auxiliary instructions for revising the previous responses. Note that the subscripts indicate turn indices within the multi-turn setting. Since we do not consider external feedback (e.g., compiler outcomes, other LMs’ feedback), is fixed for each turn.
@@ 4.2 Setup
We experiment with two prompting-based self-correction methods, Self-Refine (Madaan et al., 2023) and SETS (Chen et al., 2025)—both demonstrated exclusively on proprietary LLMs in their original work. To verify the effectiveness of SLMs, all experiments are conducted using the pre-trained Qwen2.5-1.5B (Yang et al., 2024). Self-Refine uses 3-shot prompting, whereas SETS relies on test-time scaling. Since SETS does not generate a fixed initial response, we estimate Accuracy@t1 by averaging the accuracy of 20 sampled initial outputs. Accuracy@t2 is similarly computed after all correction steps. However, because SETS verifies each sample and only applies correction to those deemed correct, improvements between initial and final responses cannot be consistently measured. Instead, we report feedback accuracy in Appendix A.
We further investigate the influence of external feedback by employing strong teacher models,222We use gpt-4.1-2025-04-14 (OpenAI, 2025) for teacher model. which we refer to as External-Refine and External-SETS. For comparison, we also introduce a simple correction prompting method that first generates an initial response with a 3-shot prompt, then prompts for correction without disclosing the correctness of the initial output. Additional details including prompts and hyperparameters are provided in Appendix G.
@@ 5.1 Multi-turn MDP
As outlined in §3, our goal is to enhance model outputs across multiple turns by leveraging the response history. Since the model’s own generated history naturally defines a sequential decision process, we formulate the process of self-correction as a multi-turn Markov decision process (Qu et al., 2024), where the policy is defined as
In this formulation, the state at each turn is denoted as , which consists of the input , the model’s prior outputs , and the prompt history . The action corresponds to generating a revised output . This action is then appended to form the next prompt for the subsequent turn. Accordingly, the policy can be rewritten as . We evaluate the correctness of the generated code using unit test execution , which assesses the functional correctness of the code against the given unit tests. The test outcome is used as the reward:
where denotes the reward for .
Under this formulation, the learning objective is to optimize a policy that maximizes the expected cumulative reward over a refinement trajectory:
where is the Kullback-Leibler (KL) coefficient that controls the strength of regularization against the reference policy .
@@ 5.2 Observed limitations
where is a binary reward, taking a value of for all-pass and for fail. The second KL-divergence term represents the regularization, and . However, we find that introducing the second KL-divergence term () leads to training collapse for SLMs, as shown in Figure 1 where the reward converges to zero.
@@ 5.3 CoCoS
Since our setting is also formulated as a multi-turn MDP, we adopt a REINFORCE-style policy gradient method (Ahmadian et al., 2024) to optimize the policy, which is widely used in single-turn RLHF (). In order to adapt to a multi-turn setting, we introduce two key modifications to the reward function. First, we introduce an accumulated reward function that incorporates a discount factor to encourage the model to consider the assessment of the previous turn; in the case of SCoRe, since it computes the reward independently at each turn. Unlike conventional RL, where discounts future rewards, we adapt its usage to amplify the contribution of the most recent turn. The accumulated reward function is defined as follows (with in our setup):
@@
The effectiveness of the progressive reward compared to the binary scheme is presented in §8.2. Accordingly, our training objective is defined as follows:
where denotes the KL coefficient. Further experimental details can be found in Appendix C.
@@
| Method | Accuracy@t1 | Accuracy@t2 | (t1, t2) | (t1, t2) | (t1, t2) |
| Boost model | 9.0% | 10.2% | 1.2% | 0.0% | 1.2% |
| Turn-SFT | 44.2% | 44.4% | 1.4% | 1.2% | 0.2% |
| Self-Corrector | 43.8% | 48.8% | 11.2% | 6.2% | 5.0% |
| ReVISE | 34.6% | 41.2% | 11.3% | 4.9% | 6.4% |
| Turn-RL | 46.4% | 48.6% | 4.0% | 1.8% | 2.2% |
| CoCoS | 45.0% | 54.2% | 11.0% | 1.8% | 9.2% |
@@ 6.2 Datasets
While supervised fine-tuning (SFT) allows models to learn proper code formatting through input-output pairs, online-RL lacks such structural guidance. Therefore, we perform SFT on the KodCode (Xu et al., 2025) dataset to produce parseable code so that RL training can be conducted with unit test execution. KodCode is curated to exclude data from MBPP, HumanEval, and ODEX. We measure the model’s accuracy on MBPP at both the pre-SFT and post-SFT, which we denote as the Pre-trained and Boost model, respectively, and report the results in Table 2. As shown in Table 2, we emphasize that the purpose of SFT on KodCode is not to enhance performance on the target dataset. The Boost model is the backbone for both CoCoS and all baseline comparisons.
@@ 7 Main Results
As shown in Table 3, CoCoS demonstrates the highest correction performance across all turns. Although the baselines are trained to actively revise prior responses, the metric indicates that they frequently make unnecessary revisions to already correct outputs. CoCoS, on the other hand, corrects selectively: it achieves a low rate of 1.8%, avoiding unnecessary changes, while maintaining a high rate of 11%. This results in a of 9.2%, the highest among all baselines.
@@ 8 Analyses
In this section, we analyze the discount factor in the accumulated reward function (§8.1), compare progressive and binary reward schemes (§8.2), examine multi-turn correction beyond two turns (§8.3), evaluate the model’s robustness to varied correction instructions during inference (§8.4), and conduct case studies on representative examples (§8.5).
@@ 8.4 Auxiliary instruction
To evaluate the generalization ability of our method across different prompts, we intentionally use a different instruction prompt at test time than the one used during training (as defined in Appendix G.1). Instead of the fixed training prompt, we manually rephrase it into five alternative versions for the correction task. The results are shown in Table 6, and the rephrased instructions are provided in Appendix H. While the fixed prompt used during training yields slightly better overall performance, the model exhibits no substantial degradation under instruction variation. The average performance drop at the second attempt is only 0.8%, indicating robustness to diverse prompt variations.
8.5 Case Study
To analyze the types of mistakes that CoCoS makes and how it corrects them, we conduct three case studies. The observed error types include: (1) naming errors, (2) complex logic errors, and (3) incorrect usage of code libraries. In the main text, we focus on the first category, while detailed discussions of the latter two are provided in the Appendix F.
Code 1 highlights an issue related to function name generation. Given the test cases, the model is expected to generate a function name that allows the code to execute correctly. However, in some instances, the model fails to produce the appropriate function name. In such cases, even if the core logic is correctly implemented, the test cases cannot be executed, and the result is considered a generation failure. In the following turn, when prompted to correct the previous error, the model is able to preserve the correct logic and revise only the function name, thereby resolving the issue.
@@ 9 Conclusion
Acknowledgments
This work was supported by the IITP(Institute of Information & Coummunications Technology Planning & Evaluation)-ITRC(Information Technology Research Center) grant funded by the Korea government(Ministry of Science and ICT)(IITP-2025-RS-2024-00437866, 47.5%). This work was supported by Smart HealthCare Program funded by the Korean National Police Agency(KNPA) (No. RS-2022-PT000186, 47.5%). This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH), 5%).
@@ Acknowledgments
Online reinforcement learning incurs a cost at every action step, as the model learns by interacting with the environment in real-time. Due to infrastructural constraints, we limited CoCoS training to only two turns. In the turn-wise analysis section, CoCoS consistently demonstrated accuracy gains as the number of turns increased. These results suggest that increasing the number of training turns has the potential to yield further improvements. Therefore, future work should extend our approach to support multi-turn training beyond two turns, potentially incorporating more cost-efficient strategies such as offline reinforcement learning.
For the same reason, our experiments were conducted only on a 1B-scale small language model. While we demonstrated that our methodologies are effective on small models, our approach is not limited to this setting. Moreover, evaluating its scalability to models larger than 1B parameters remains an interesting aspect to observe in future work.
@@
Appendix A Additional Experiment on SETS
@@ Appendix B Baseline Details
We generate training data using pre-trained models according to the format specified in §G.1. For each problem, we sample 10 candidate code solutions and determine their correctness based on the results of corresponding unit test cases. We then organize this data into two separate datasets: 1-turn and 2-turn. We combine the 1-turn and 2-turn datasets in a 1:1 ratio to match the data scale of other baseline methods, allowing the model to optimize initial code generation and correction capabilities.
@@ Appendix C Experimental Details
C.1 Policy Optimization
We adopt the policy gradient method using the REINFORCE Leave-One-Out (RLOO) estimator (Ahmadian et al., 2024). RLOO provides a simple and efficient baseline, and can further reduce variance when multiple online samples are available by using each sample’s reward as a baseline for others and averaging gradient estimates. We extend this approach to the multi-turn setting to train CoCoS. Our training objective, as defined in Equation 4, is as follows:
We then optimize the expected return using policy gradients with the RLOO estimator:
Here, denotes the leave-one-out baseline-adjusted reward computed over the entire trajectory. For a sampled trajectory , the trajectory-level reward is defined as:
where is the scalar reward associated with the model’s response at turn . Given such sampled trajectories, we define the leave-one-out adjusted reward as:
This estimator leverages the rewards of the other samples as a baseline, reducing variance while maintaining unbiasedness. Accordingly, the full policy gradient is computed as:
C.2 Hyperparameter
@@
| RLOO samples () | - | 2 |
C.3 Dataset
@@ Appendix E Additional SCoRe Experiment on Another Model
Appendix F Additional Case Studies
We present two additional case studies that illustrate distinct types of errors: (2) complex logic errors and (3) incorrect usage of code libraries. These additional cases serve to complement the discussion in the main text by illustrating error types that extend beyond naming issues. They underscore CoCoS’s capacity to detect and remediate a diverse set of programming errors.
Code 2 illustrates an error involving the coordination of multiple functions. The model initially fails to implement the correct interaction logic across functions, resulting in an incorrect solution. In the subsequent turn, it is able to revise the necessary parts while preserving the correct subcomponents, effectively resolving the issue.
Lastly, Code 3 demonstrates a misuse of a Python method. The model generates an invalid call to str.translate() that is not supported in the current version of Python. In the subsequent turn, it corrects the implementation by replacing the unsupported method with a valid character filtering loop.
Appendix G Prompts
In this section, we present the prompts used for prompt-based approaches and for CoCoS training and evaluation. The prompts for implementing prompting-based approaches are described in §G.2 and §G.3. Finally, the prompts used for CoCoS training and evaluation are reported in §G.1.
G.1 Main experiments